# 3D-RAD: A Comprehensive 3D Radiology Med-VQA Dataset with Multi-Temporal Analysis and Diverse Diagnostic Tasks

### v1

@@ Abstract
Medical Visual Question Answering (Med-VQA) holds significant potential for clinical decision support, yet existing efforts primarily focus on 2D imaging with limited task diversity. This paper presents 3D-RAD, a large-scale dataset designed to advance 3D Med-VQA using radiology CT scans. The 3D-RAD dataset encompasses six diverse VQA tasks: anomaly detection, image observation, medical computation, existence detection, static temporal diagnosis, and longitudinal temporal diagnosis. It supports both open- and closed-ended questions while introducing complex reasoning challenges, including computational tasks and multi-stage temporal analysis, to enable comprehensive benchmarking. Extensive evaluations demonstrate that existing vision-language models (VLMs), especially medical VLMs exhibit limited generalization, particularly in multi-temporal tasks, underscoring the challenges of real-world 3D diagnostic reasoning. To drive future advancements, we release a high-quality training set 3D-RAD-T of 136,195 expert-aligned samples, showing that fine-tuning on this dataset could significantly enhance model performance. Our dataset and code, aiming to catalyze multimodal medical AI research and establish a robust foundation for 3D medical visual understanding, are publicly available at https://github.com/Tang-xiaoxiao/M3D-RAD.
@@ 1 Introduction
In this paper, we introduce 3D-RAD, a large-scale dataset designed for effective training and comprehensive benchmarking of 3D Med-VQA tasks using radiology CT and text. The 3D-RAD consists of 34k data entries across six challenging VQA task types—spanning open-ended (e.g., anomaly detection, image observation, medical computation) and closed-ended (e.g., existence detection, static temporal diagnosis, longitudinal temporal diagnosis). It covers over 18 diseases and supports advanced reasoning tasks, such as medical computation and temporal analysis, making it highly suitable for practical evaluation (see Figure 2). The dataset adheres to data usage regulations and ethical considerations. We leverage strong large language models (LLMs) for annotation, with rigorous consistency checks across LLMs to ensure high-quality annotations, supplemented by extensive human validation for quality control. Additionally, we provide a training subset, 3D-RAD-T, with fine-tuned models demonstrating substantial performance gains. By releasing 3D-RAD, we aim to provide a reliable resource to advance 3D Med-VQA, facilitating research into network architectures and pretraining strategies optimized for 3D multimodal clinical reasoning. Our main contributions are summarized as follows:
Figure 1: Qualitative Comparison with Different Types of Dataset.
@@
Table 1: Comparison of 3D-RAD with existing Med-VQA datasets.
Dataset Modality Dataset Scale Tasks Covered Temporal Reasoning 3D Support Quality Check
VQA-RAD [11] 2D (X-ray, CT) 315 images 3,515 QA Modality, anatomy, abnormalities Dual Annotation
SLAKE [13] 2D 642 images 14,028 QA Function, anatomy -
PathVQA [14] 2D (Pathology) 4,998 images 32,799 QA Histopathology Manual Validation
VQA-Med [12] 2D $\sim$5K images Classification, description Manual Validation
M3D-VQA [16] 3D (CT) 120K QA Slice finding, spatial reasoning, diagnosis -
3D-RAD (Ours) 3D (CT) 16,188 images, 136K QA across 6 tasks Detection, quantification, diagnosis, progression LLM Score + Human Validation + Consistency Check
@@ 2 Related Work
To bridge this gap, we introduce 3D-RAD, a large-scale and clinically grounded 3D Med-VQA benchmark. It comprises 16,188 CT cases with diagnostic reports and defines six task types spanning recognition, detection, diagnosis, anatomical reasoning, and multi-temporal progression. Unlike prior datasets focused on static or single-task settings, 3D-RAD supports dynamic, longitudinal reasoning by incorporating follow-up imaging and historical reports. It provides high-quality annotations, expert-validated QA, and systematic evaluation across multiple VLMs. Extensive experiments demonstrate current models struggle with complex 3D and temporal tasks, especially multi-temporal reasoning. 3D-RAD thus offers a scalable, extensible platform to advance multimodal reasoning, fairness, and clinical applicability in 3D Med-VQA.
Figure 2: Definition of 3D-RAD Open-End and Closed-End Tasks
@@
The dataset construction process involves a semi-automated pipeline combining radiology report parsing, expert annotation, and temporal case matching. We further incorporate both closed-ended and open-ended questions to capture different forms of clinical reasoning, and ensure label quality through multi-temporal validation. The final dataset includes 34k QA samples, each paired with 3D CT scans, offering a diverse and challenging benchmark for advancing 3D Med-VQA research.
Figure 3: Dataset Construction Pipeline
@@ 3.1 Tasks Definitions
Source Category Task 2 Task 3 Task 4 Task 5 Task 6 Task 7 Overall
3D-RAD-Bench Images 980 1,858 656 1,304 169 169 2,662
Patients 651 1,008 331 1,304 169 169 1,304
Q-A Pairs 1,024 2,666 1,002 23,472 2,873 2,873 33,910
3D-RAD-T Images 2,045 5,670 1,148 5,565 774 774 13,526
Patients 1,697 4,443 587 5,565 774 774 9,951
Q-A Pairs 2,081 6,055 1,573 100,170 13,158 13,158 136,195
Table 2: Image, Patient, and QA Distribution across 3D-RAD Training and Benchmark Sets
@@ 3.2.1 Dataset Construction
Figure 4: A Concise Prompt Example of Task 1-2
Open-Ended Questions (Task 1-3): To construct data for open-ended tasks, we extract the Finding and Impression fields from the CT-RATE clinic report text. For each task, we design a corresponding prompt. Figure 4 presents a concise example of the prompt, with the comprehensive prompt detailed in the  Figure 9, Figure 10, Figure 11.
For Tasks 1–2, we construct QA pairs by injecting image-associated clinical report text into carefully designed prompts and feeding them into GPT-4o-mini[37]. For Task 3, we adopt a two-stage pipeline. In the first stage, we extract sentences containing quantitative measurements (e.g., "5 mm", "3×5 mm") from the clinical reports. In the second stage, these extracted sentences are inserted into prompts and input to GPT-4o-mini to generate QA pairs. To ensure the answers remain concise and aligned with the QA format—rather than drifting toward explanatory responses—we constrain generated answers to approximately five words. After the initial generation, all QA pairs undergo a post-hoc quality control process, including automatic scoring and manual sampling-based validation, resulting in a curated high-quality dataset.
Figure 5: Data Distribution of 3D-RAD Dataset
@@
We construct closed-ended question-answer pairs by selecting samples with multi-label abnormality annotations from the CT-RATE dataset. For Task 4, we design 18 binary classification prompts, each aligned with one of the predefined abnormality labels. To promote question diversity, the prompt pool is shuffled and assigned sequentially across labels. Answers are directly obtained from the annotations, with “0” mapped to No and “1” to Yes. For Task 5, we focus on patients with two or more scans. For each diagnostic label, we analyze the historical labels from the previous n–1 scans and classify the current status into one of four temporal categories. Task 6 adopts the same answer construction strategy as Task 5, deriving responses based on findings from the previous n–1 scans. For question generation, we design three modular prompt components: (1) temporal label templates, (2) interpretation templates, and (3) base question prompts. These components are independently shuffled and concatenated to promote diversity. Temporal labels are generated by mapping historical scan results (1 for presence, 0 for absence) and embedded into the question. Full template details are provided in the  Figure 12, Figure 13, Figure 14.
Data Validation:
To ensure the quality of 3D-RAD, we develop a semi-automated scoring and filtering pipeline. Each generated QA pair is evaluated using a GPT-based framework across five dimensions—Visual Verifiability, Specificity & Clarity, Answer Appropriateness, Q-A Alignment, and Linguistic Quality—rated on a 1–5 scale. Pairs scoring below 3 in any dimension or with an average score below 3 are discarded. The remaining QA pairs are ranked by average score, and to encourage diversity, only the top 10 instances are retained for each unique question or answer. High-quality samples are then selected top-down to construct the final dataset (see Figure 4). To validate the reliability of our scoring model (GPT-4o-mini), we sample 600 QA pairs and compare its ratings with those of DeepSeek-R1 [38], LLaMA3.3-70B, and LLaMA3-8B [39] using a shared evaluation rubric. As shown in Figure 8 and Figure 8, GPT-4o-mini exhibits the highest agreement with other models on high-quality QA pairs (measured by Top-100 Overlap and NDCG@100), despite minor variation on lower-scoring samples due to scoring conservativeness. Since only high-scoring QA pairs are retained, GPT-4o-mini is adopted as the primary scoring model.
@@ 3.3 Data Statistics
3D-RAD consists of two subsets: a high-quality benchmark set and a large-scale training set, 3D-RAD-T. Based on the 1,304 validation cases and 20,000 training cases in the original CT-RATE dataset, we initially generate over 400k QA pairs using our task-specific templates. To ensure quality and reduce redundancy, we perform automatic deduplication, scoring, and filtering. This results in 33,910 QA pairs for the benchmark set and 136,195 for the final training set. A detailed breakdown is provided in Table 2. To assess the linguistic and thematic diversity of 3D-RAD, we conduct a word cloud analysis across tasks (in the  Figure 27) [40]. The question distributions vary notably by tasks, reflecting distinct focuses while maintaining coherence within each task’s clinical objective. This demonstrates the richness and task-alignment of our constructed QA corpus. To analyze the biomedical terminology covered in our benchmark dataset, we utilize the en_ner_bionlp13cg_md model[41] to extract named entities belonging to the categories PATHOLOGICAL_FORMATION, TISSUE, ORGANISM_SUBDIVISION, and ORGAN. Based on the extracted entities, we generate a unified word cloud to visualize the vocabulary distribution. Figure 5 demonstrates that our dataset encompasses a wide range of disease-related and biomedical terms, highlighting its linguistic diversity and domain coverage.
Table 3: Finetuned Results of M3D-RAD (Llama2-7B) and M3D-RAD (Phi3-4B) on All Tasks
@@
Anomaly Detection BLEU 10.69 31.28 ↑+20.59 16.31 39.66↑+23.35
@@
Image Observation BLEU 9.10 25.25↑+16.15 15.06 33.28↑+18.22
Rouge 18.64 33.76↑+15.12 23.19 42.45↑+19.26
BERTScore 86.07 89.16↑+3.09 87.11 90.72↑+3.61
Medical Computation BLEU 15.95 30.54↑+14.59 2.55 20.85↑+18.30
Rouge 23.24 36.06↑+12.82 5.63 30.38↑+24.75
BERTScore 91.50 94.65↑+3.15 85.74 93.25↑+7.51
@@ 4.1 Experimental Setup
Anomaly Detection BLEU 13.48 10.69 16.31 16.42
@@ 4.2 Experiment Reults
M3D-RAD Model. To assess the utility of 3D-RAD, we fine-tune two M3D model variants with different parameter scales, thereby constructing the M3D-RAD models. Table 3 reports the average performance across all tasks before and after fine-tuning (see  Figure 22, Figure 23, Figure 24, Figure 25, Figure 26 for detailed results). As shown, both small and large variants of M3D-RAD benefit consistently from fine-tuning across all tasks, confirming the utility of our dataset. Notably, for the newly introduced Task 5 and Task 6—both involving multi-phase reasoning—the models initially performed poorly, with accuracies around 20%. After fine-tuning with our dataset, their accuracies increased significantly, reaching over 70%, highlighting the large room for improvement in current models when dealing with temporally structured tasks. Nevertheless, despite this substantial gain, the performance on these new tasks remains lower than that on Task 5, a traditional single-phase task, suggesting that while temporal cues help, they do not fully close the gap in reasoning complexity.
@@
Figure 6: Finetuned results of models on Task 1–3
Zero-Shot Evaluation on Existing Models. We conducted zero-shot evaluation of several state-of-the-art 3D medical vision-language models on our benchmark to assess their generalization capabilities in Table 4 (see Appendix for subtask results). On conventional single-phase tasks (Task 1, Task 2, and Task 4), M3D-4B and OmniV exhibited the strongest performance. Interestingly, M3D-7B achieved superior results on Task 3, which requires numerical reasoning, suggesting stronger arithmetic capabilities. In contrast, RadFM achieved the best results on our newly proposed tasks, indicating better generalization to temporally structured QA formats. These results suggest that current models exhibit task-specific strengths, but none demonstrate universally strong performance across all clinical tasks.This task-specific variation underscores the need for more comprehensive benchmarks that span diverse clinical reasoning types—including computation, visualization, and multi-phase inference—as covered in our dataset. The complete experimental results on all subtasks are illustrated in  Figure 17, Figure 18, Figure 19, Figure 20, Figure 21.

### v2

@@ Abstract
Medical Visual Question Answering (Med-VQA) holds significant potential for clinical decision support, yet existing efforts primarily focus on 2D imaging with limited task diversity. This paper presents 3D-RAD, a large-scale dataset designed to advance 3D Med-VQA using radiology CT scans. The 3D-RAD dataset encompasses six diverse VQA tasks: anomaly detection, image observation, medical computation, existence detection, static temporal diagnosis, and longitudinal temporal diagnosis. It supports both open- and closed-ended questions while introducing complex reasoning challenges, including computational tasks and multi-stage temporal analysis, to enable comprehensive benchmarking. Extensive evaluations demonstrate that existing vision-language models (VLMs), especially medical VLMs exhibit limited generalization, particularly in multi-temporal tasks, underscoring the challenges of real-world 3D diagnostic reasoning. To drive future advancements, we release a high-quality training set 3D-RAD-T of 136,195 expert-aligned samples, showing that fine-tuning on this dataset could significantly enhance model performance. Our dataset and code, aiming to catalyze multimodal medical AI research and establish a robust foundation for 3D medical visual understanding, are publicly available.
@@ 1 Introduction
In this paper, we introduce 3D-RAD, a large-scale dataset designed for effective training and comprehensive benchmarking of 3D Med-VQA tasks using radiology CT and text. The 3D-RAD consists of 170K data entries across six challenging VQA task types—spanning open-ended (e.g., anomaly detection, image observation, medical computation) and closed-ended (e.g., existence detection, static temporal diagnosis, longitudinal temporal diagnosis). It covers over 18 diseases and supports advanced reasoning tasks, such as medical computation and temporal analysis, making it highly suitable for practical evaluation (see Figure 2). The dataset adheres to data usage regulations and ethical considerations. We leverage strong large language models (LLMs) for annotation, with rigorous consistency checks across LLMs to ensure high-quality annotations, supplemented by extensive human validation for quality control. Additionally, we provide a training subset, 3D-RAD-T, with fine-tuned models demonstrating substantial performance gains. By releasing 3D-RAD, we aim to provide a reliable resource to advance 3D Med-VQA, facilitating research into network architectures and pretraining strategies optimized for 3D multimodal clinical reasoning. Our main contributions are summarized as follows:
Figure 1: Qualitative Comparison Across Dataset Types. Prior work (e.g., VQA-RAD; top) focuses on 2D open- and closed-ended VQA, whereas our 3D-RAD (bottom) additionally includes 3D imaging and multi-temporal tasks.
@@
Table 1: Comparison of 3D-RAD with Existing Med-VQA Datasets across Modality, Scale, Tasks Covered, Temporal Reasoning, 3D Support, and Quality Check.
Dataset Modality
Dataset
Scale
Tasks
Covered
Temporal
Reasoning
3D
Support
Quality
Check
VQA-RAD [11] 2D (X-ray, CT)
315 images
3,515 QA
Modality, anatomy,
abnormalities
Dual Annotation
SLAKE [13] 2D
642 images
14,028 QA
Function, anatomy
PathVQA [14] 2D (Pathology)
4,998 images
32,799 QA
Histopathology
Manual Validation
VQA-Med [12] 2D $\sim$5K images
Classification,
description
Manual Validation
M3D-VQA [16] 3D (CT) 120K QA
Slice finding,
spatial reasoning,
diagnosis
3D-RAD (Ours) 3D (CT)
16,188 images,
170K QA
across 6 tasks
Detection, quantification,
diagnosis, progression
LLM Score + Human
Validation + Consistency Check
@@ 2 Related Work
To bridge this gap, we introduce 3D-RAD, a large-scale and clinically grounded 3D Med-VQA benchmark. It defines six task types spanning recognition, detection, diagnosis, anatomical reasoning, and multi-temporal progression. Unlike prior datasets focused on static or single-task settings, 3D-RAD supports dynamic, longitudinal reasoning by incorporating follow-up imaging and historical reports. It provides high-quality annotations, expert-validated QA, and systematic evaluation across multiple VLMs. Extensive experiments demonstrate current models struggle with complex 3D and temporal tasks, especially multi-temporal reasoning. 3D-RAD thus offers a scalable, extensible platform to advance multimodal reasoning, fairness, and clinical applicability in 3D Med-VQA.
Figure 2: Definitions of Open-Ended and Closed-Ended Tasks in 3D-RAD. Different colors indicate distinct tasks; items sharing the same color represent different subtasks within that task.
@@
The dataset construction process involves a semi-automated pipeline combining radiology report parsing, expert annotation, and temporal case matching. We further incorporate both closed-ended and open-ended questions to capture different forms of clinical reasoning, and ensure label quality through multi-temporal validation. The final dataset comprises samples, each paired with 3D CT scans, providing a diverse and challenging benchmark for advancing 3D Med-VQA research. We partition the dataset into two distinct subsets to ensure strict separation between training and evaluation, preventing data leakage and enabling fair, reproducible comparisons.
  • 3D-RAD-Bench: approximately 34K QA pairs with 2,662 images across six representative tasks, designed for standardized evaluation)
  • 3D-RAD-T (Train): approximately 136K QA pairs with 13,526 images constructed independently from the benchmark using the same procedure, with no overlap, expansion, or resampling.
@@ 3.1 Tasks Definitions
Figure 3: 3D-RAD Dataset Construction Pipeline. Left: meta dataset with a 3D scan, clinical report, and structured labels. Middle: QA construction—open-ended (Tasks 1–2) from selected report sentences with a five-dimension quality check; numeric (Task 3) from measurements; closed-ended (Tasks 4–6) via prompt templates and choice lists, with temporal indices for Task 6. Right: representative QA examples for Tasks 1–6.
Figure 4: A Concise Prompt Example for Tasks 1–2. See Figure 11 for the full prompt.
Figure 5: Data Distribution of 3D-RAD Dataset. Left: Organ distribution derived from named‑entity recognition (NER) over all 3D‑RAD QA pairs, showing coverage of major thoracic and upper‑abdominal structures. Right: QA‑based NER term statistics—bar chart of the most frequent entities/concepts alongside a word cloud that highlights the broader long‑tail vocabulary of anatomical and pathological descriptors.
@@ 3.2.1 Dataset Construction
Open-Ended Questions (Task 1-3): To construct data for open-ended tasks, we extract the Finding and Impression fields from the CT-RATE clinic report text. For each task, we design a corresponding prompt. Figure 4 presents a concise example of the prompt, with the comprehensive prompt detailed in Figure 11, Figure 12, and Figure 13.
For Tasks 1–2, we construct QA pairs by injecting image-associated clinical report text into carefully designed prompts and feeding them into GPT-4o-mini[39]. For Task 3, we adopt a two-stage pipeline. In the first stage, we extract sentences containing quantitative measurements (e.g., "5 mm", "3×5 mm") from the clinical reports. In the second stage, these extracted sentences are inserted into prompts and input to GPT-4o-mini to generate QA pairs.
To diversify question forms and reduce repetitive patterns, we adopt the “6W” framework [40, 41] and, following expert validation, select three clinically common starters—“what,” “where,” and “which”—for our templates. For each report, each starter is used twice to produce a list of six that is then shuffled and assigned, ensuring diversity across starters, image content, and textual cues; this strategy promotes uniqueness and balanced distribution. To mitigate redundancy, we truncate high-frequency questions or answers by retaining only the first 10 instances when an item appears more than 10 times, preserving essential semantic and contextual variety while better reflecting common clinical inquiries than extreme deduplication. We also conduct entity coverage analysis using named entity recognition to assess the distribution of anatomical entities in questions; summary statistics are reported in Figure 5, with task-wise word clouds in Figure 29 further demonstrating organ and expression diversity. To ensure the answers remain concise and aligned with the QA format—rather than drifting toward explanatory responses—we constrain generated answers to approximately five words. After the initial generation, all QA pairs undergo a post-hoc quality control process, including automatic scoring and manual sampling-based validation, resulting in a curated high-quality dataset.
Table 2: Distributions of Images, Patients, and QA pairs across tasks in the 3D‑RAD training and benchmark sets.
Source Category Task 1 Task 2 Task 3 Task 4 Task 5 Task 6 Overall
3D-RAD-Bench Images 1,858 980 656 1,304 169 169 2,662
Patients 1,008 651 331 1,304 169 169 1,304
Q-A Pairs 2,666 1,024 1,002 23,472 2,873 2,873 33,910
3D-RAD-T (Train) Images 5,670 2,045 1,148 5,565 774 774 13,526
Patients 4,443 1,697 587 5,565 774 774 9,951
Q-A Pairs 6,055 2,081 1,573 100,170 13,158 13,158 136,195
@@
We construct closed-ended question-answer pairs by selecting samples with multi-label abnormality annotations from the CT-RATE dataset. For Task 4, we design 18 binary classification prompts, each aligned with one of the predefined abnormality labels. To promote question diversity, the prompt pool is shuffled and assigned sequentially across labels. Answers are directly obtained from the annotations, with “0” mapped to No and “1” to Yes. For Task 5, we focus on patients with two or more scans. For each diagnostic label, we analyze the historical labels from the previous n–1 scans and classify the current status into one of four temporal categories. Task 6 adopts the same answer construction strategy as Task 5, deriving responses based on findings from the previous n–1 scans. For question generation, we design three modular prompt components: (1) temporal label templates, (2) interpretation templates, and (3) base question prompts. These components are independently shuffled and concatenated to promote diversity. Temporal labels are generated by mapping historical scan results (1 for presence, 0 for absence) and embedded into the question. Full template details are provided in Figure 14, Figure 15, and Figure 16.
Data Validation: To ensure the quality of 3D-RAD, we develop a semi-automated scoring and filtering pipeline. Each generated QA pair is evaluated using a GPT-based framework across five dimensions—Visual Verifiability, Specificity & Clarity, Answer Appropriateness, Q-A Alignment, and Linguistic Quality—rated on a 1–5 scale. Pairs scoring below 3 in any dimension or with an average score below 3 are discarded. The remaining QA pairs are ranked by average score, and to encourage diversity, only the top 10 instances are retained for each unique question or answer. High-quality samples are then selected top-down to construct the final dataset (see Figure 4). To validate the reliability of our scoring model (GPT-4o-mini), we sample 600 QA pairs and compare its ratings with those of DeepSeek-R1 [42], LLaMA3.3-70B, and LLaMA3-8B [43] using a shared evaluation rubric. As shown in Figure 8 and Figure 8, GPT-4o-mini exhibits the highest agreement with other models on high-quality QA pairs (measured by Top-100 Overlap and NDCG@100), despite minor variation on lower-scoring samples due to scoring conservativeness. Since only high-scoring QA pairs are retained, GPT-4o-mini is adopted as the primary scoring model.
@@ 3.3 Data Statistics
3D-RAD consists of two subsets: a high-quality benchmark set and a large-scale training set, 3D-RAD-T. Based on the 1,304 validation cases and 20,000 training cases in the original CT-RATE dataset, we initially generate over 400k QA pairs using our task-specific templates. To ensure quality and reduce redundancy, we perform automatic deduplication, scoring, and filtering. This results in 33,910 QA pairs for the benchmark set and 136,195 for the final training set. A detailed breakdown is provided in Table 2. To assess the linguistic and thematic diversity of 3D-RAD, we conduct a word cloud analysis across tasks (in Figure 29) [44]. The question distributions vary notably by tasks, reflecting distinct focuses while maintaining coherence within each task’s clinical objective. This demonstrates the richness and task-alignment of our constructed QA corpus. To analyze the biomedical terminology covered in our benchmark dataset, we utilize the en_ner_bionlp13cg_md model[45] to extract named entities belonging to the categories PATHOLOGICAL_FORMATION, TISSUE, ORGANISM_SUBDIVISION, and ORGAN. Based on the extracted entities, we generate a unified word cloud to visualize the vocabulary distribution. Figure 5 demonstrates that our dataset encompasses a wide range of disease-related and biomedical terms, highlighting its linguistic diversity and domain coverage.
Table 3: Finetuned Results of M3D-RAD (Llama2-7B) and M3D-RAD (Phi3-4B) on All Tasks. The best performance is highlighted in bold, while the second-best is underlined. Red numbers denote the performance gain over the base model after fine-tuning.
@@
Anomaly Detection BLEU 9.10 25.25↑+16.15 15.06 33.28↑+18.22
Rouge 18.64 33.76↑+15.12 23.19 42.45↑+19.26
BERTScore 86.07 89.16↑+3.09 87.11 90.72↑+3.61
Image Observation BLEU 10.69 31.28 ↑+20.59 16.31 39.66↑+23.35
@@
Medical Computation BLEU 15.95 30.54↑+14.59 2.55 33.52↑+30.97
Rouge 23.24 36.06↑+12.82 5.63 36.46↑+30.83
BERTScore 91.50 94.65↑+3.15 85.74 94.86↑+9.12
@@ 4.1 Experimental Setup
Anomaly Detection BLEU 11.00 9.10 15.06 13.47
Rouge 17.62 18.64 23.19 25.72
BERTScore 86.76 86.07 87.11 88.21
Image Observation BLEU 13.48 10.69 16.31 16.42
@@ 4.2 Experiment Reults
M3D-RAD Model. To assess the utility of 3D-RAD, we fine-tune two M3D model variants with different parameter scales, thereby constructing the M3D-RAD models. Table 3 reports the average performance across all tasks before and after fine-tuning (see Figure 18, Figure 19, Figure 20, Figure 21, and Figure 22 for detailed results). As shown, both small and large variants of M3D-RAD benefit consistently from fine-tuning across all tasks, confirming the utility of our dataset. Notably, for the newly introduced Task 5 and Task 6—both involving multi-phase reasoning—the models initially performed poorly, with accuracies around 20%. After fine-tuning with our dataset, their accuracies increased significantly, reaching over 70%, highlighting the large room for improvement in current models when dealing with temporally structured tasks. Nevertheless, despite this substantial gain, the performance on these new tasks remains lower than that on Task 5, a traditional single-phase task, suggesting that while temporal cues help, they do not fully close the gap in reasoning complexity.
@@
Figure 6: Finetuned Results of Models on Task 1–3. Background shading denotes each task, while bar colors indicate different proportions of the training data used for fine‑tuning.
Zero-Shot Evaluation on Existing Models. We conducted zero-shot evaluation of several state-of-the-art 3D medical vision-language models on our benchmark to assess their generalization capabilities in Table 4 (see Appendix for subtask results). On conventional single-phase tasks (Task 1, Task 2, and Task 4), M3D-4B and OmniV exhibited the strongest performance. Interestingly, M3D-7B achieved superior results on Task 3, which requires numerical reasoning, suggesting stronger arithmetic capabilities. In contrast, RadFM achieved the best results on our newly proposed tasks, indicating better generalization to temporally structured QA formats. These results suggest that current models exhibit task-specific strengths, but none demonstrate universally strong performance across all clinical tasks.This task-specific variation underscores the need for more comprehensive benchmarks that span diverse clinical reasoning types—including computation, visualization, and multi-phase inference—as covered in our dataset. The complete experimental results on all subtasks are illustrated in  Figure 23, Figure 24, Figure 25, Figure 26, and Figure 27.